The costs of materials consumed in producing good units in the Production Department of Jacobs Company were $38,000 and $39,125 for June and July, respectively. The number of equivalent units produced in June and July was 4,000 and 4,250, respectively. Which of the following best describes the change in the cost of materials between the two months?

a.The cost of materials increased by $0.28 per unit, indicating an unfavorable change

b.The cost of materials decreased by $0.29 per unit, indicating an improvement.

c.The cost of materials increased by $0.88 per unit, indicating an unfavorable change.

d.The cost of materials increased by $1,125, indicating an unfavorable change.
Answer #1

Cost per equivalent unit of June = 38000/4000 = 9.50

Cost per equivalent unit of July = 39125/4250 = 9.21

So cost per equivalent unit decrease by (9.50-9.21) = 0.29

So answer is b) b.The cost of materials decreased by $0.29 per unit, indicating an improvement.